OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution of the equation. -12x-5y=-54 answers; (-3,18) (-2,22) (-1,15) (-2,17)

OpenStudy (anonymous):

@baby1234567 @douglaswinslowcooper @tkhunny @rose_101 @rose21

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

(-1,15)

OpenStudy (texaschic101):

-12x - 5y = -54 -12(-1) - 5(15) = -54 12 - 75 = -54 well..it is not (-1,15) sub in your answer choices and see which one satisfies the equation. Lets try (-3,18) -12 (-3) - 5(18) = -54 36 - 90 = -54 -54 = -54 (correct) answer is : (-3,18)

OpenStudy (anonymous):

^ thank you!

OpenStudy (texaschic101):

no problem :)

OpenStudy (jojo4eva):

sorry my internet was actin up but if u have any others juss lemme kno

OpenStudy (anonymous):

okay, I will. (: